# GRAVITY for MATISSE

### v1

@@ Abstract
Context. MATISSE, the mid-infrared spectro-imaging instrument of VLTI, was designed to deliver its advertised performance when paired with an external second generation fringe tracker. Science observation started in 2019, demonstrating imaging capabilities and faint science target observations. Now, The GRAVITY fringe tracker stabilizes the MATISSE fringes which allows using all spectroscopic modes and improves sensitivity and data accuracy.
Aims. We present how the MATISSE and GRAVITY instruments were adapted to make the GRAVITY fringe tracker work with MATISSE, under the umbrella of the aptly-named GRA4MAT project, led by ESO in collaboration with the two instrument consortia.
Methods. We detail the software modifications needed to implement an acquisition and observing sequence specific to GRA4MAT, including simultaneous fringe tracking and chopping and a narrow off-axis capability inspired by the galactic center and exoplanet capability of GRAVITY. We explain the modified data collection and reduction processes. We show how we leveraged the recent fringe tracker upgrade to implement features specific to its use with MATISSE, e.g. fringe jumps mitigation with an improved group delay control and simultaneous fringe tracking and chopping with a new state machine.
Results. We successfully demonstrate significant improvements to the MATISSE instrument. Observations can now be performed at higher spectral resolutions of up to and across the full LM bands at once. Long detector integration times, made possible with stabilized fringes, have improved the LM-bands sensitivity by a factor of 10. Low flux biases in coherently-reduced N-band data have been eliminated. The L-band transfer function is now higher and more stable. We finally illustrate the scientific potential of GRA4MAT with a preview of the first exoplanet observation made by MATISSE on Pictoris b.
@@
A second-generation fringe tracker for the Very Large Telescope Interferometer (VLTI) was originally foreseen to support the second-generation instruments GRAVITY (GRAVITY Collaboration et al., 2017) and MATISSE (Lopez et al., 2022). A call for studies for such an instrument was issued by ESO in late 2008. In a context where available resources were being allocated in priority to the second generation instruments, including the former PRIMA astrometry project (Sahlmann et al., 2013), none of the proposed solutions (Blind et al., 2010; Meisner et al., 2012) materialized. Whereas GRAVITY came with an internal fringe tracker (Menu et al., 2012; Choquet et al., 2014; Lacour et al., 2019), MATISSE came without the fringe tracker it needed to achieve its advertised performance and its circumstellar disks and active galactic nuclei science objectives. This is how the idea of using the GRAVITY fringe tracker to stabilize the optical path differences of MATISSE came to life, as an alternative to this second generation fringe tracker.
@@
GRAVITY for MATISSE, or GRA4MAT in short, was first proposed by ESO in 2013 as an alternative to the second-generation fringe tracker. The concept was supported by the MATISSE consortium and endorsed by the ESO council in 2014. Subsequently, concept studies were prepared by ESO and the MATISSE consortium in 2014-2015, ahead of a 2015 phase A review, conducted under the umbrella of the VLTI facility project (Woillez et al., 2015). As such, GRA4MAT is the last project of the so-called second generation VLTI. The project only acquired momentum in 2019, after the completion over the 2015-2016 period of the significant upgrade of the interferometer infrastructure (Woillez et al., 2015) required to host the GRAVITY and MATISSE instruments, and the installation in late 2018 of a new adaptive optics system (Woillez et al., 2019, NAOMI) on the auxiliary telescopes. The project passed final design review in August 2019, following prototyping activities carried out shortly before the review. GRA4MAT was implemented, tested, and commissioned in Paranal over the period September 2019 - March 2023, including an extended pause due to the COVID-19 reaching Europe and Chile in 2020, delays resulting from the failure of the L-band dispersion wheel inside MATISSE between January 2021 and February 2022, and needed synergies with the GRAVITY fringe tracker real-time controller upgrade (Abuter et al., 2016; Nowak et al., 2024) between April 2022 and March 2023.
@@
GRA4MAT is not the first instrument to pair a near-infrared fringe tracker and a mid-infrared instrument. On the Keck Interferometer (Colavita et al., 2013), the two-telescopes Keck Nuller (Serabyn et al., 2012) was the first instrument of this kind. It was primarily used to survey exo-zodiacal lights around nearby bright stars (Millan-Gabet et al., 2011; Mennesson et al., 2014), and developed many of the techniques further extended by the GRA4MAT project. On the Large Binocular Telescope (LBT), a similar nulling instrument was also developed (LBTI, Defrère et al., 2016) for a similar exo-zodiacal purpose (Ertel et al., 2020). On the VLTI, one of the objectives of the PRIMA project (Delplancke, 2008) was to use the FSU fringe tracker (Sahlmann et al., 2009) to support the two-telescopes MIDI instrument (Leinert et al., 2003), with the so-called MIDI+FSU (Müller, 2012; Müller et al., 2014); it however remained at the level of an experiment.
@@
GRA4MAT can also be considered as a precursor to instruments currently under development. The L-band nuller NOTT (formerly Hi-5 Defrère et al., 2018) of the Asgard visitor instrument suite proposal (Martinod et al., 2023) aims at bringing nulling technique pioneered in the US to VLTI. GRA4MAT even has connections with the mid-infrared ELT instrument METIS (Brandl et al., 2021), where water vapour dispersion introduces a flat wavefront mismatch between wavefront sensing in the visible and high contrast imaging in the thermal infrared. As such, GRA4MAT provides direct measurements of this effect on baselines lengths comparable to the ELT diameter (Absil et al., 2022). This represents an interesting parallel between adaptive optics on a diameter telescope and fringe tracking on a baseline interferometer.
@@ 2 Implementation
GRA4MAT is almost exclusively a software implementation; none of the hardware of GRAVITY, MATISSE, or VLTI was modified. It however benefited from the upgrade of the GRAVITY fringe tracker, which also included a hardware component (Abuter et al., 2016; Nowak et al., 2024).
@@ 2.1 Hardware implementation
The layout of GRA4MAT is the layout of MATISSE and GRAVITY in the VLTI laboratory, illustrated in Fig. 1. The GRA4MAT setup uses the same MATISSE feeding optics that split the LMN bands for MATISSE and the K band for IRIS, the infrared tip/tilt sensor of VLTI. This time, the KH bands are sent to GRAVITY where K band is used for fringe tracking, and H band for field guiding. The simple long-wavelength-reflecting and short-wavelength-transmissive dichroic plate, dictated by the relative position of the two instruments, could not be manufactured for technical feasibility reasons. Instead, fold mirrors are added to operate MATISSE in transmission and GRAVITY or IRIS (Gitton et al., 2004) in reflection. The dichroic and fold mirrors are inserted in the beam only when MATISSE is operating in standalone or GRA4MAT modes; to switch from MATISSE standalone to GRA4MAT, the GRAVITY Feeding Optics are inserted in the beam.
@@ 2.2 Software implementation
The software architecture of GRA4MAT considers GRAVITY as a sub-system of MATISSE. When the MATISSE instrument operates in GRA4MAT mode, its state also depends on the state of GRAVITY. The MATISSE acquisition and observation sequencers directly control GRAVITY, without going through the Interferometer Supervisory Software, the interface to the VLTI infrastructure for all VLTI instruments. The functions available from the GRAVITY instrument and fringe tracker workstations are accessed directly by MATISSE; whereas functions only available from GRAVITY local control units are accessed through a gateway running on the GRAVITY workstation. Feedback from GRAVITY is obtained through live database links from the GRAVITY workstation to the MATISSE workstation.
@@ 2.3 Acquisition sequence
Rather than re-implementing the GRAVITY acquisition procedure in the MATISSE instrument, which would have required access from MATISSE to many low-level GRAVITY functions, MATISSE prepares an acquisition Observing Block (OB) following the single off-axis template and remotely triggers its execution on GRAVITY in the so-called “Bob_Stephanie”111Bob is the Broker of Observations Block, well known to the VLT users. Usually, one single Bob runs per instrument. In the GRA4MAT case, since another Bob runs on the GRAVITY side, we gave it another name: Stephanie.. This OB skips sending target coordinates to an already tracking VLTI but configures GRAVITY for single off-axis operations to send all the K-band light to the fringe tracker. This represents a transmission gain of x2 over the on-axis mode usually employed by GRAVITY for single target observations. This choice optimizes the fringe tracker limiting magnitude but prevents any light from reaching the GRAVITY spectrometer.
@@
The acquisition sequence for GRAVITY follows its standard procedure. The acquisition camera field loop adjusts the internal tip-tilt mirror of the instrument to couple the target light into the single mode fibres of the fringe tracker. A small synchronous modulation of the same internal mirror, and demodulation of the coupled flux, removes any pointing mismatch between the single mode fibres and the acquisition camera. Searching and tracking fringes completes the GRAVITY acquisition procedure. Then the acquisition on MATISSE resumes.
@@
First, MATISSE switches the actuator used by GRAVITY for field control from the internal Tip Tilt Piston (TTP) actuator to the common-path field steering mirrors of the telescope Star Separators (STS, Nijenhuis & Giesen, 2005). This makes GRAVITY stabilize the field for both instruments. MATISSE first offsets the internal tip-tilt actuator of GRAVITY to a reference value that corresponds to the co-alignment between the two instruments. This offset is compensated by the common-path field steering mirror of the Star Separator (STS) to keep the target position unchanged on the fringe tracker fibres. Then MATISSE performs a field acquisition measurement and sends combined offsets to the internal tip/tilt mirrors of GRAVITY and the STSs to refine the co-alignment between the two instruments. After this operation, since the GRAVITY field derotator remains static in single off-axis mode, the two instruments remain co-aligned all through the night and this acquisition does not have to be repeated after the first object of the night.
@@
Second, MATISSE switches the actuator used by GRAVITY for piston control from the internal tip/tilt to the common-path VLTI delay lines. This switch comes with an adjustment of the actuator model used by the predictive Kalman controller of the fringe tracker. Indeed, the VLTI delay lines have a longer latency (see Fig. 3). Again, this makes GRAVITY stabilize the piston for both instruments. MATISSE completes the acquisition sequence with an offset to its internal delay lines to match the optical path differences of GRAVITY.
@@ 2.4 Coherent flux observations
The GRA4MAT observation sequence remains identical to the one with MATISSE standalone. After a sky measurement, non-chopped fringes are recorded for the requested number of Beam Commuting Device (BCD) cycles (see section 2.4 of Lopez et al., 2022).
Except for the piston actuator replaced by the main delay lines, the fringe tracker operates for MATISSE the same way it does for GRAVITY. The only critical aspect is related to how GRA4MAT configures the fringe tracker for the management of dispersion and the mitigation of fringe jumps. Since its upgrade (Nowak et al., 2024), the fringe tracker is capable of periodically updating the set point of its group delay control loop. A regular update of this set point is critical to follow the slowly evolving dispersion, and keeping fringe jumps under control. GRA4MAT uses this feature to schedule the updates between the MATISSE Detector Integration Times, the same way GRAVITY standalone does it between its own DITs. When the L-band DITs are shorter than , GRA4MAT schedules the updates between the OPD modulation cycles applied by MATISSE on the L-band side of the instrument. When the L-band DITs are longer than 1 second, the updates are scheduled before each L-band DIT. This is done because the group delay set point is constant and not able to follow the evolving dispersion over hundreds of seconds; its validity in time is therefore limited. For long L-band DITs, the fringe tracker will therefore not keep the phase stable for the duration of a MATISSE Optical Path Difference (OPD) cycle, and coherent integrations by the MATISSE pipeline must consider the group delay tracked by GRAVITY.
@@ 2.5 Narrow off-axis offsets
The narrow off-axis mode of GRA4MAT was not part of the initial objectives of the project. It was added along the way, in light of the pioneering galactic center (e.g. GRAVITY Collaboration et al., 2018) and exoplanet (e.g. GRAVITY Collaboration et al., 2019) results obtained with GRAVITY. With this mode, MATISSE is now able to observe any faint target within a VLTI field of view ( and diameter with Unit Telescopes and Auxilliary Telescopes, respectively) centred on a bright fringe tracking target.
To perform a narrow off-axis observation, the user provides in the observing template a sequence of cumulative on-sky offsets starting from the fringe tracker target. This allows to sequentially observe multiple possibly fainter targets around the fringe tracking target serving as a calibrator. Such a sequence is illustrated in Table 1. The requested relative offsets are applied in tip/tilt and piston, by the MATISSE observation sequencer, at the start of each BCD cycle.
@@
In tip/tilt, this offset is a combined offset between the common-path field steering mirrors of the STSs and the GRAVITY-only internal tip/tilt mirrors. As a result, the MATISSE pinhole moves to the requested offset while the GRAVITY fringe tracker remains in place. It should be noted that this offset strategy does not continuously compensate for field rotation. Whereas this is not an issue with the ATs that deliver a stable field orientation to the interferometric laboratory, this affects the autonomy for the UTs and their rotating fields. Therefore, a recalculation of the narrow off-axis offset is performed before each BCD state, leaving the UTs field rotation uncompensated over no more than exposure intervals.

### v2

@@ Abstract
Context. MATISSE, the mid-infrared spectro-imaging instrument of VLTI, was designed to deliver its designed performance when paired with an external second-generation fringe tracker. Science observations started in 2019, demonstrating imaging capabilities and faint science target observations. Now, the GRAVITY fringe tracker stabilises the MATISSE fringes, which allows the use of all spectroscopic modes and improves sensitivity and data accuracy.
Aims. We present how the MATISSE and GRAVITY instruments were adapted to make the GRAVITY fringe tracker work with MATISSE, under the umbrella of the aptly named GRA4MAT project, led by ESO in collaboration with the two instrument consortia.
Methods. We detail the software modifications needed to implement an acquisition and observing sequence specific to GRA4MAT, including simultaneous fringe tracking and chopping and a narrow off-axis capability inspired by the Galactic Centre and exoplanet capability of GRAVITY. We explain the modified data collection and reduction processes. We show how we leveraged the recent fringe tracker upgrade to implement features specific to its use with MATISSE, for example mitigation of fringe jumps with an improved group delay control, and simultaneous fringe tracking and chopping with a new state machine.
Results. We successfully demonstrate significant improvements to the MATISSE instrument. Observations can now be performed at higher spectral resolutions of up to and across the full LM bands at once. Long detector integration times, made possible with stabilised fringes, have improved the LM-band sensitivity by a factor of 10. Low flux biases in coherently reduced N-band data have been eliminated. The L-band transfer function is now higher and more stable. We finally illustrate the scientific potential of GRA4MAT with a preview of the first exoplanet observation made by MATISSE on $β$ Pictoris b.
@@
A second-generation fringe tracker for the Very Large Telescope Interferometer (VLTI) was originally foreseen to support the second-generation instruments GRAVITY (GRAVITY Collaboration et al., 2017) and MATISSE (Lopez et al., 2022). A call for studies for such an instrument was issued by ESO in late 2008. In a context where available resources were being allocated in priority to the second-generation instruments, including the former PRIMA astrometry project (Sahlmann et al., 2013), none of the proposed solutions (Blind et al., 2010; Meisner et al., 2012) materialised. Whereas GRAVITY came with an internal fringe tracker (Menu et al., 2012; Choquet et al., 2014; Lacour et al., 2019), MATISSE came without the fringe tracker it needed to achieve its planned performance and its circumstellar disks and active galactic nuclei science objectives. This is how the idea of using the GRAVITY fringe tracker to stabilise the optical path differences of MATISSE came to life, as an alternative to this second-generation fringe tracker.
@@
GRAVITY for MATISSE, hereafter GRA4MAT, was first proposed by ESO in 2013 as an alternative to the second-generation fringe tracker. The concept was supported by the MATISSE consortium and endorsed by the ESO council in 2014. Subsequently, concept studies were prepared by ESO and the MATISSE consortium in 2014-2015, ahead of a 2015 phase A review, conducted under the umbrella of the VLTI facility project (Woillez et al., 2015). As such, GRA4MAT is the last project of the second-generation VLTI. The project only acquired momentum in 2019, after the completion over the 2015-2016 period of the significant upgrade of the interferometer infrastructure (Woillez et al., 2015) required to host the GRAVITY and MATISSE instruments, and the installation in late 2018 of a new adaptive optics system (Woillez et al., 2019, NAOMI) on the auxiliary telescopes. The project passed final design review in August 2019, following prototyping activities carried out shortly before the review. GRA4MAT was implemented, tested, and commissioned in Paranal over the period September 2019–March 2023, including an extended pause due to the COVID-19 pandemic reaching Europe and Chile in 2020, delays resulting from the failure of the L-band dispersion wheel inside MATISSE between January 2021 and February 2022, and needed synergies with the GRAVITY fringe tracker real-time controller upgrade (Abuter et al., 2016; Nowak et al., 2024) between April 2022 and March 2023.
@@
GRA4MAT is not the first instrument to pair a near-infrared fringe tracker and a mid-infrared instrument. On the Keck Interferometer (Colavita et al., 2013), the two-telescope Keck Nuller (Serabyn et al., 2012) was the first instrument of this kind. It was primarily used to survey exo-zodiacal lights around nearby bright stars (Millan-Gabet et al., 2011; Mennesson et al., 2014), and developed many of the techniques further extended by the GRA4MAT project. On the Large Binocular Telescope (LBT), a similar nulling instrument (LBTI) was also developed (Defrère et al., 2016) for a similar exo-zodiacal purpose (Ertel et al., 2020). On the VLTI, one of the objectives of the PRIMA project (Delplancke, 2008) was to use the FSU fringe tracker (Sahlmann et al., 2009) to support the two-telescope MIDI instrument (Leinert et al., 2003), with a setup called MIDI+FSU (Müller, 2012; Müller et al., 2014); however, it remained at the level of an experiment.
@@
GRA4MAT can also be considered as a precursor to instruments currently under development. The L-band nuller NOTT (formerly Hi-5 Defrère et al., 2018) of the Asgard visitor instrument suite proposal (Martinod et al., 2023) aims at bringing the nulling technique pioneered in the US to VLTI. GRA4MAT even has connections with the mid-infrared ELT instrument METIS (Brandl et al., 2021), where water vapour dispersion introduces a flat wavefront mismatch between wavefront sensing in the visible and high-contrast imaging in the thermal infrared. As such, GRA4MAT provides direct measurements of this effect on baseline lengths comparable to the ELT diameter (Absil et al., 2022). This represents an interesting parallel between adaptive optics on a diameter telescope and fringe tracking on a baseline interferometer.
@@ 2 Implementation
GRA4MAT is almost exclusively a software implementation; none of the hardware of GRAVITY, MATISSE, or VLTI was modified. However, it benefited from the upgrade of the GRAVITY fringe tracker, which also included a hardware component (Abuter et al., 2016; Nowak et al., 2024).
@@ 2.1 Hardware implementation
The layout of GRA4MAT is the layout of MATISSE and GRAVITY in the VLTI laboratory, illustrated in Fig. 1. The GRA4MAT set-up uses the same MATISSE feeding optics that split the LMN bands for MATISSE and the K band for IRIS, the infrared tip-tilt sensor of VLTI. This time, the KH bands are sent to GRAVITY where the K band is used for fringe tracking and the H band for field guiding. The simple long-wavelength-reflecting and short-wavelength-transmissive dichroic plate, dictated by the relative position of the two instruments, could not be manufactured for technical feasibility reasons. Instead, fold mirrors are used to operate MATISSE in transmission and GRAVITY or IRIS (Gitton et al., 2004) in reflection. The dichroic and fold mirrors are inserted in the beam only when MATISSE is operating in standalone or GRA4MAT modes; to switch from MATISSE standalone to GRA4MAT, the GRAVITY feeding optics are inserted in the beam.
@@ 2.2 Software implementation
The software architecture of GRA4MAT considers GRAVITY as a subsystem of MATISSE. When the MATISSE instrument operates in GRA4MAT mode, its state also depends on the state of GRAVITY. The MATISSE acquisition and observation sequencers directly control GRAVITY without going through the Interferometer Supervisory Software, the interface to the VLTI infrastructure for all VLTI instruments. The functions available from the GRAVITY instrument and fringe tracker workstations are accessed directly by MATISSE, whereas functions only available from GRAVITY local control units are accessed through a gateway running on the GRAVITY workstation. Feedback from GRAVITY is obtained through live database links from the GRAVITY workstation to the MATISSE workstation.
@@ 2.3 Acquisition sequence
Rather than re-implementing the GRAVITY acquisition procedure in the MATISSE instrument, which would require access from MATISSE to many low-level GRAVITY functions, MATISSE prepares an acquisition observing block (OB) following the single off-axis template and remotely triggers its execution on GRAVITY using the tool called Bob_Stephanie.111‘Bob’ is the Broker of Observations Block, well known to VLT users. Usually, one single Bob runs per instrument. Since another Bob runs on the GRAVITY side, in the GRA4MAT case we gave it another name: Stephanie. This OB skips sending target coordinates to an already tracking VLTI, but configures GRAVITY for single off-axis operations to send all the K-band light to the fringe tracker. This represents a two-fold transmission gain over the on-axis mode usually employed by GRAVITY for single-target observations. This choice optimises the fringe tracker limiting magnitude, but prevents any light from reaching the GRAVITY spectrometer.
@@
The acquisition sequence for GRAVITY follows the standard procedure. The acquisition camera field loop adjusts the internal tip-tilt mirror of the instrument to couple the target light into the single-mode fibres of the fringe tracker. A small synchronous modulation of the same internal mirror, and demodulation of the coupled flux, removes any pointing mismatch between the single-mode fibres and the acquisition camera. Searching and tracking fringes completes the GRAVITY acquisition procedure. Then the acquisition on MATISSE resumes.
@@
First, MATISSE switches the actuator used by GRAVITY for field control from the internal tip-tilt-piston (TTP) actuator to the common-path field steering mirrors of the telescope star separators (STS, Nijenhuis & Giesen, 2005). This makes GRAVITY stabilise the field for both instruments. MATISSE first offsets the internal tip-tilt actuator of GRAVITY to a reference value that corresponds to the co-alignment between the two instruments. This offset is compensated by the common-path field steering mirror of the star separator (STS) to keep the target position unchanged on the fringe tracker fibres. Then MATISSE performs a field acquisition measurement and sends combined offsets to the internal tip-tilt mirrors of GRAVITY and the STSs to refine the co-alignment between the two instruments. After this operation, since the GRAVITY field derotator remains static in single off-axis mode, the two instruments remain co-aligned all through the night, and this acquisition does not have to be repeated after the first object of the night.
@@
Second, MATISSE switches the actuator used by GRAVITY for piston control from the internal tip-tilt to the common-path VLTI delay lines. This switch comes with an adjustment of the actuator model used by the predictive Kalman controller of the fringe tracker. The VLTI delay lines have a longer latency (see Fig. 3). Again, this makes GRAVITY stabilise the piston for both instruments. MATISSE completes the acquisition sequence with an offset to its internal delay lines to match the optical path differences of GRAVITY.
@@ 2.4 Coherent flux observations
The GRA4MAT observation sequence remains identical to that with MATISSE standalone. After a sky measurement, non-chopped fringes are recorded for the requested number of beam commuting device (BCD) cycles (see section 2.4 in Lopez et al., 2022).
Except for the piston actuator replaced by the main delay lines, the fringe tracker operates for MATISSE the same way it does for GRAVITY. The only critical aspect is related to how GRA4MAT configures the fringe tracker for the management of dispersion and the mitigation of fringe jumps. Since its upgrade (Nowak et al., 2024), the fringe tracker is capable of periodically updating the set point of its group delay control loop. A regular update of this set point is critical to follow the slowly evolving dispersion, and keeping fringe jumps under control. GRA4MAT uses this feature to schedule the updates between the MATISSE detector integration times, the same way GRAVITY standalone does it between its own DITs. When the L-band DITs are shorter than , GRA4MAT schedules the updates between the OPD modulation cycles applied by MATISSE on the L-band side of the instrument. When the L-band DITs are longer than , the updates are scheduled before each L-band DIT. This is done because the group delay set point is constant and is not able to follow the evolving dispersion over hundreds of seconds; its validity in time is therefore limited. For long L-band DITs, the fringe tracker therefore does not keep the phase stable for the duration of a MATISSE optical path difference (OPD) cycle, and coherent integrations by the MATISSE pipeline must consider the group delay tracked by GRAVITY.
@@ 2.5 Narrow off-axis offsets
The narrow off-axis mode of GRA4MAT was not part of the initial objectives of the project. It was added along the way in light of the pioneering Galactic Centre results (e.g. GRAVITY Collaboration et al., 2018) and exoplanet results (e.g. GRAVITY Collaboration et al., 2019) obtained with GRAVITY. With this mode, MATISSE is now able to observe any faint target within a VLTI field of view ( and diameter with Unit Telescopes and Auxilliary Telescopes, respectively) centred on a bright fringe tracking target.
To perform a narrow off-axis observation, the user provides in the observing template a sequence of cumulative on-sky offsets starting from the fringe tracker target. This allows multiple possibly fainter targets to be sequentially observed, around the fringe tracking target serving as a calibrator. Such a sequence is illustrated in Table 1. The requested relative offsets are applied in tip-tilt and piston, by the MATISSE observation sequencer, at the start of each BCD cycle.
@@
In tip-tilt, this offset is a combined offset between the common-path field steering mirrors of the STSs and the GRAVITY-only internal tip-tilt mirrors. As a result, the MATISSE pinhole moves to the requested offset while the GRAVITY fringe tracker remains in place. It should be noted that this offset strategy does not continuously compensate for field rotation. Although this is not an issue with the ATs that deliver a stable field orientation to the interferometric laboratory, this affects the autonomy for the UTs and their rotating fields. Therefore, a recalculation of the narrow off-axis offset is performed before each BCD state, leaving the UTs field rotation uncompensated over no more than exposure intervals.